OpenStudy (anonymous):

what is 3^square root 54

OpenStudy (anonymous):

Thanks man!

OpenStudy (whpalmer4):

Is that supposed to be \[3^\sqrt{54}\] or \[\sqrt[3]{54}\]

OpenStudy (whpalmer4):

My guess is the latter...let's assume it is. If we factor 54 into its prime components, we get 54 = 2*27 = 2*3*9 = 2*3*3*3 so \[\sqrt[3]{54} = \sqrt[3]{2*3*3*3}\]Remembering that if you have a square root, you can take out one of every pair of identical factors, what do you think you might do here with a cube root? Hint: \[\sqrt[3]{8} = \sqrt[3]{2*2*2} = 2\]